### v2

@@ Abstract
We prove that the distribution of the product of two correlated normal random variables with non-zero means and arbitrary variances is infinitely divisible. We also obtain exact formulas for the probability density function of the sum of independent copies of such random variables.
@@
AMS 2010 Subject Classification: Primary 60E05; 60E07; 62E15; Secondary 33C15; 60E10
@@ 1 Introduction
where $K_{\nu}(x)$ is a modified Bessel function of the second kind (defined in Appendix A). If one of the means is equal to zero and $\rho=0$, then the PDF (1) reduces to a single infinite series (see [9, 27]). Without loss of generality, set $\mu_{Y}=0$. Then the density (1) reduces to
$f_{Z}(x)=\frac{1}{\pi\sigma_{X}\sigma_{Y}}\exp\bigg{(}\!-\frac{\mu_{X}^{2}}{2\sigma_{X}^{2}}\bigg{)}\sum_{j=0}^{\infty}\frac{\mu_{X}^{2j}|x|^{j}}{(2j)!\sigma_{X}^{3k}\sigma_{Y}^{j}}K_{j}\bigg{(}\frac{|x|}{\sigma_{X}\sigma_{Y}}\bigg{)},\quad x\in\mathbb{R}.$ (1.2)
We further observe that if $\mu_{X}/\sigma_{X}=\rho\mu_{Y}/\sigma_{Y}$ or $\mu_{Y}/\sigma_{Y}=\rho\mu_{X}/\sigma_{X}$, then the formula (1) simplifies to a single infinite series. If $\mu_{X}/\sigma_{X}=\rho\mu_{Y}/\sigma_{Y}$, then, for $x\in\mathbb{R}$,
@@
If $\mu_{Y}/\sigma_{Y}=\rho\mu_{X}/\sigma_{X}$, then, for $x\in\mathbb{R}$,
$f_{Z}(x)$ $=\frac{1}{\pi\sigma_{X}\sigma_{Y}\sqrt{1-\rho^{2}}}\exp\bigg{\{}-\frac{\mu_{X}^{2}}{2\sigma_{X}^{2}}+\frac{\rho x}{\sigma_{X}\sigma_{Y}(1-\rho^{2})}\bigg{\}}$
×$\quad\times\sum_{j=0}^{\infty}\frac{1}{(2j)!}\bigg{(}\frac{\mu_{X}}{\sigma_{X}}\bigg{)}^{2j}\bigg{(}\frac{|x|}{\sigma_{X}\sigma_{Y}}\bigg{)}^{j}K_{j}\bigg{(}\frac{|x|}{\sigma_{X}\sigma_{Y}(1-\rho^{2})}\bigg{)}.$ (1.4)
@@
$f_{Z}(x)=\frac{1}{\pi\sigma_{X}\sigma_{Y}\sqrt{1-\rho^{2}}}\exp\bigg{(}\frac{\rho x}{\sigma_{X}\sigma_{Y}(1-\rho^{2})}\bigg{)}K_{0}\bigg{(}\frac{|x|}{\sigma_{X}\sigma_{Y}(1-\rho^{2})}\bigg{)},\quad x\in\mathbb{R}.$ (1.5)
In the case $\mu_{X}=\mu_{Y}=0$, the formula (1.5) for the PDF of $Z$ generalises in a natural manner to yield the following exact formula for the PDF of $\overline{Z}_{n}=n^{-1}\sum_{i=1}^{n}Z_{i}$, which was derived independently by [22, 23, 32]:
@@
In this paper, we obtain exact formulas for the PDF of the sum $S_{n}=\sum_{i=1}^{n}Z_{i}$ for the full range of parameter values; exact formulas for the PDF of the sample mean $\overline{Z}_{n}$ follow from a simple rescaling, as noted in Remark 2.3. We provide an infinite series representation of the PDF (Theorem 2.1) and an integral representation of the PDF (Theorem 2.8). As is the case for the the PDF of the product $Z$, our series representation of the PDF of the sum $S_{n}$ simplifies for certain parameter values, and we note these simplifications in Remark 2.5 and Theorem 2.6.
@@ 2.1 Infinite series representations of the density
Let $\mu_{X},\mu_{Y}\in\mathbb{R}$, $\sigma_{X},\sigma_{Y}>0$ and $-1<\rho<1$. Then, for $x\in\mathbb{R}$,
$f_{S_{n}}(x)$ $=\frac{(1-\rho^{2})^{n/2-1}}{2^{n-1}\sigma_{X}\sigma_{Y}}\exp\bigg{(}-{\color[rgb]{0,0,0}\frac{n}{2(1-\rho^{2})}\bigg{(}\frac{\mu_{X}^{2}}{\sigma_{X}^{2}}+\frac{\mu_{Y}^{2}}{\sigma_{Y}^{2}}-\frac{2\rho\mu_{X}\mu_{Y}}{\sigma_{X}\sigma_{Y}}\bigg{)}}-\frac{|x|}{\sigma_{X}\sigma_{Y}(1+\rho\,\mathrm{sgn}(x))}\bigg{)}$
×$\quad\times\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/{\color[rgb]{0,0,0}8})^{k}}{k!\Gamma(n/2+a_{j,k}(x))}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\bigg{)}^{j}\bigg{(}\frac{\mu_{X}}{\sigma_{X}}-\frac{\mu_{Y}}{\sigma_{Y}}\bigg{)}^{2j}\bigg{(}\frac{1-\rho}{1+\rho}\bigg{)}^{k-j}$
@@
Suppose $\mu_{X},\mu_{Y}\in\mathbb{R}$, $\sigma_{X},\sigma_{Y}>0$ and $-1<\rho<1$. Then, the PDF $f_{S_{n}}(x)$ is bounded for all $x\in\mathbb{R}$ if and only if $n=1$.
@@
It was shown in [14] that the PDF of $Z=_{d}S_{1}$ is unimodal with a logaritmic singularity at the mode $x=0$ (this can also be inferred from formula (2.7) by using the limiting forms (A.34) and (A.35)). For $n\geq 2$, the PDF (2.7) of $S_{n}$ is, however, bounded for all $x\in\mathbb{R}$, as can be seen from applying the limiting form to (A.35) to determine the behaviour as $x\rightarrow 0$, and due to the fact that the function $U(a,b,x)$ only has (possible) singularities in the limits $x\rightarrow 0$ and $x\rightarrow\pm\infty$, and the PDF is clearly bounded as $x\rightarrow\pm\infty$ (a property satisfied by any PDF). ∎
@@
$f_{\overline{Z}_{n}}(x)$ $=\frac{n(1-\rho^{2})^{n/2-1}}{2^{n-1}\sigma_{X}\sigma_{Y}}\exp\bigg{(}-{\color[rgb]{0,0,0}\frac{n}{2(1-\rho^{2})}\bigg{(}\frac{\mu_{X}^{2}}{\sigma_{X}^{2}}+\frac{\mu_{Y}^{2}}{\sigma_{Y}^{2}}-\frac{2\rho\mu_{X}\mu_{Y}}{\sigma_{X}\sigma_{Y}}\bigg{)}}-\frac{n|x|}{\sigma_{X}\sigma_{Y}(1+\rho\,\mathrm{sgn}(x))}\bigg{)}$
×$\quad\times\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/{\color[rgb]{0,0,0}8})^{k}}{k!\Gamma(n/2+a_{j,k}(x))}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\bigg{)}^{j}\bigg{(}\frac{\mu_{X}}{\sigma_{X}}-\frac{\mu_{Y}}{\sigma_{Y}}\bigg{)}^{2j}\bigg{(}\frac{1-\rho}{1+\rho}\bigg{)}^{k-j}$
@@
$f_{S_{n}}(x)$ $=\frac{(1-\rho^{2})^{n/2-1}}{2^{n-1}\sigma_{X}\sigma_{Y}}\exp\bigg{(}-\frac{n\mu_{X}^{2}}{\sigma_{X}^{2}(1+\rho)}-\frac{|x|}{\sigma_{X}\sigma_{Y}(1+\rho\,\mathrm{sgn}(x))}\bigg{)}\sum_{k=0}^{\infty}\frac{{\color[rgb]{0,0,0}(n/2)}^{k}}{k!\Gamma(n/2+a_{0,k}(x))}$
@@
$f_{S_{n}}(x)$ $=\frac{(1-\rho^{2})^{n/2-1}}{2^{n-1}\sigma_{X}\sigma_{Y}}\exp\bigg{(}-\frac{n\mu_{X}^{2}}{\sigma_{X}^{2}(1-\rho)}-\frac{|x|}{\sigma_{X}\sigma_{Y}(1+\rho\,\mathrm{sgn}(x))}\bigg{)}\sum_{k=0}^{\infty}\frac{{\color[rgb]{0,0,0}(n/2)}^{k}}{k!\Gamma(n/2+a_{k,k}(x))}$
@@
We therefore see that formula (2.7) reduces to a single infinite series in the cases $\mu_{X}/\sigma_{X}=\mu_{Y}/\sigma_{Y}$ and $\mu_{X}/\sigma_{X}=-\mu_{Y}/\sigma_{Y}$. This is in contrast to the formula (1) of [9] that reduces to a single infinite series if $\mu_{X}/\sigma_{X}=\rho\mu_{Y}/\sigma_{Y}$ or $\mu_{Y}/\sigma_{Y}=\rho\mu_{X}/\sigma_{X}$ (see (1.3) and (1.4)).
@@
Suppose $\rho=0$, $\mu_{Y}=0$, and $\mu_{X}\in\mathbb{R}$ and $\sigma_{X},\sigma_{Y}>0$. Then, for $x\in\mathbb{R}$,
@@
Theorem 2.6 extends in a natural way the PDF $f_{Z}(x)$ described in (1.2) to the PDF $f_{S_{n}}(x)$. On calculating $(2j)!=\Gamma(2j+1)$ using the gamma duplication formula $\Gamma(2x)=\pi^{-1/2}2^{2x-1}\Gamma(x)\Gamma(x+1/2)$ see ([24, Section 5.5(iii)]) we see that formula (2.8) with $n=1$ indeed reduces to (1.2).
@@
$\frac{{\color[rgb]{0,0,0}-}(\mu_{X}^{2}+\mu_{Y}^{2}-2\rho\mu_{X}\mu_{Y})t^{2}+2\mu_{X}\mu_{Y}\mathrm{i}t}{2[1-(1+\rho)\mathrm{i}t][1+(1-\rho)\mathrm{i}t]}$
${\color[rgb]{0,0,0}=-\frac{(\mu_{X}^{2}+\mu_{Y}^{2}-2\rho\mu_{X}\mu_{Y})}{2(1-\rho^{2})}+\frac{1}{4(1-\rho^{2})}\bigg{(}\frac{(1+\rho)(\mu_{X}-\mu_{Y})^{2}}{(1+(1-\rho)\mathrm{i}t)}+\frac{(1-\rho)(\mu_{X}+\mu_{Y})^{2}}{(1-(1+\rho)\mathrm{i}t)}\bigg{)}}.$
@@
×$\quad\times\exp\bigg{(}{\color[rgb]{0,0,0}\frac{n}{4(1-\rho^{2})}\bigg{(}\frac{(1+\rho)(\mu_{X}-\mu_{Y})^{2}}{(1+(1-\rho)\mathrm{i}t)}+\frac{(1-\rho)(\mu_{X}+\mu_{Y})^{2}}{(1-(1+\rho)\mathrm{i}t)}\bigg{)}}\bigg{)}$
@@
×$\quad\times\sum_{k=0}^{\infty}\frac{(n/{\color[rgb]{0,0,0}4})^{k}}{k!{\color[rgb]{0,0,0}(1-\rho^{2})^{k}}}\bigg{(}\frac{(1+\rho)(\mu_{X}-\mu_{Y})^{2}}{(1+(1-\rho)\mathrm{i}t)}+\frac{(1-\rho)(\mu_{X}+\mu_{Y})^{2}}{(1-(1+\rho)\mathrm{i}t)}\bigg{)}^{k}$
$=\exp\bigg{(}-\frac{n(\mu_{X}^{2}+\mu_{Y}^{2}-2\rho\mu_{X}\mu_{Y})}{2(1-\rho^{2})}\bigg{)}\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/{\color[rgb]{0,0,0}4})^{k}}{k!{\color[rgb]{0,0,0}(1-\rho^{2})^{k}}}\binom{k}{j}(1+\rho)^{j}(\mu_{X}-\mu_{Y})^{2j}$
@@
$f_{S_{n}}(x)$ $=\frac{1}{2\pi}\exp\bigg{(}-\frac{n(\mu_{X}^{2}+\mu_{Y}^{2}-2\rho\mu_{X}\mu_{Y})}{2(1-\rho^{2})}\bigg{)}\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/{\color[rgb]{0,0,0}4})^{k}}{k!{\color[rgb]{0,0,0}(1-\rho^{2})^{k}}}\binom{k}{j}(1+\rho)^{j}(\mu_{X}-\mu_{Y})^{2j}$
@@
$=\frac{1}{2\pi}\exp\bigg{(}-\frac{n(\mu_{X}^{2}+\mu_{Y}^{2}-2\rho\mu_{X}\mu_{Y})}{2(1-\rho^{2})}\bigg{)}\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/{\color[rgb]{0,0,0}4})^{k}}{k!(1-\rho^{2})^{n/2+k}}\binom{k}{j}$
@@
$f_{S_{n}}(x)$ $=\frac{1}{2\pi}\exp\bigg{(}-\frac{n(\mu_{X}^{2}+\mu_{Y}^{2}-2\rho\mu_{X}\mu_{Y})}{2(1-\rho^{2})}\bigg{)}\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/{\color[rgb]{0,0,0}4})^{k}}{k!(1-\rho^{2})^{n/2+k}}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\bigg{)}^{j}$
@@
×$\quad\times\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/{\color[rgb]{0,0,0}8})^{k}}{k!\Gamma(n/2+k-j)}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\bigg{)}^{j}(\mu_{X}-\mu_{Y})^{2j}\bigg{(}\frac{1-\rho}{1+\rho}\bigg{)}^{k-j}$
@@
$\quad\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/{\color[rgb]{0,0,0}8})^{k}}{k!\Gamma(n/2+j)}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\bigg{)}^{j}(\mu_{X}-\mu_{Y})^{2j}\bigg{(}\frac{1-\rho}{1+\rho}\bigg{)}^{k-j}$
@@
$=\exp\bigg{(}-\frac{n\mu^{2}}{2}\bigg{)}\sum_{k=0}^{\infty}\frac{1}{k!}\bigg{(}\frac{n\mu^{2}}{2}\bigg{)}^{j}\frac{1}{(1+t^{2})^{n/2+k}}.$
@@
$f_{S_{n}}(x)$ $=\frac{1}{2\pi}\exp\bigg{(}-\frac{n\mu^{2}}{2}\bigg{)}\sum_{k=0}^{\infty}\frac{1}{k!}\bigg{(}\frac{n\mu^{2}}{2}\bigg{)}^{j}\int_{-\infty}^{\infty}\frac{\mathrm{e}^{-\mathrm{i}xt}}{(1+t^{2})^{n/2+k}}\,\mathrm{d}t$

### v3

@@ Abstract
We prove that the distribution of the product of two correlated normal random variables with arbitrary means and arbitrary variances is infinitely divisible. We also obtain exact formulas for the probability density function of the sum of independent copies of such random variables.
@@
AMS 2020 Subject Classification: Primary 60E05; 60E07; 62E15; Secondary 33C15; 60E10
@@ 1 Introduction
where $K_{\nu}(x)$ is a modified Bessel function of the second kind. In Appendix A, we define the modified Bessel function of the second kind and the other special functions used in this paper. We also state some relevant basic properties of these special functions that are required in this paper; we will indeed draw heavily on special function theory in this paper.
If one of the means, say $\mu_{Y}$, is equal to zero and $\rho=0$, then the PDF (1) reduces to a single infinite series (see [9, 28]):
$f_{Z}(x)=\frac{1}{\pi\sigma_{X}\sigma_{Y}}\exp\bigg{(}\!-\frac{\mu_{X}^{2}}{2\sigma_{X}^{2}}\bigg{)}\sum_{j=0}^{\infty}\frac{\mu_{X}^{2j}|x|^{j}}{(2j)!\sigma_{X}^{3j}\sigma_{Y}^{j}}K_{j}\bigg{(}\frac{|x|}{\sigma_{X}\sigma_{Y}}\bigg{)},\quad x\in\mathbb{R}.$ (1.2)
We further observe that if $\mu_{X}/\sigma_{X}=\rho\mu_{Y}/\sigma_{Y}$ or $\mu_{Y}/\sigma_{Y}=\rho\mu_{X}/\sigma_{X}$, then the formula (1) simplifies to a single infinite series. For example, if $\mu_{X}/\sigma_{X}=\rho\mu_{Y}/\sigma_{Y}$, then, for $x\in\mathbb{R}$,
@@
$f_{Z}(x)=\frac{1}{\pi\sigma_{X}\sigma_{Y}\sqrt{1-\rho^{2}}}\exp\bigg{(}\frac{\rho x}{\sigma_{X}\sigma_{Y}(1-\rho^{2})}\bigg{)}K_{0}\bigg{(}\frac{|x|}{\sigma_{X}\sigma_{Y}(1-\rho^{2})}\bigg{)},\quad x\in\mathbb{R},$ (1.4)
and formula (1.4) generalises in a natural manner to yield the following exact formula for the PDF of $\overline{Z}_{n}=n^{-1}\sum_{i=1}^{n}Z_{i}$, which was derived independently by [23, 24, 33]:
@@
In this paper, we obtain exact formulas for the PDF of the sum $S_{n}=\sum_{i=1}^{n}Z_{i}$ for the full range of parameter values; exact formulas for the PDF of the sample mean $\overline{Z}_{n}$ follow from a simple rescaling, as noted in Remark 2.3. We provide an infinite series representation of the PDF (Theorem 2.1) and an integral representation of the PDF (Theorem 2.8). As is the case for the PDF of the product $Z$, our series representation of the PDF of the sum $S_{n}$ simplifies for certain parameter values, and we note these simplifications in Remark 2.5 and Theorem 2.6. The integral representation of the PDF is rather useful and has recently been applied by [15] to derive the entire asymptotic expansion for the PDF and tail probabilities of the sum $S_{n}$ for the full range of parameter values.
@@ 2.1 Infinite series representations of the density
For $x\in\mathbb{R}$,
$f_{S_{n}}(x)$ $=\frac{(1-\rho^{2})^{n/2-1}}{2^{n-1}\sigma_{X}\sigma_{Y}}\exp\bigg{(}-\frac{n}{2(1-\rho^{2})}\bigg{(}\frac{\mu_{X}^{2}}{\sigma_{X}^{2}}+\frac{\mu_{Y}^{2}}{\sigma_{Y}^{2}}-\frac{2\rho\mu_{X}\mu_{Y}}{\sigma_{X}\sigma_{Y}}\bigg{)}-\frac{|x|}{\sigma_{X}\sigma_{Y}(1+\rho\,\mathrm{sgn}(x))}\bigg{)}$
×$\quad\times\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/8)^{k}}{k!\Gamma(n/2+a_{j,k}(x))}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\bigg{)}^{j}\bigg{(}\frac{\mu_{X}}{\sigma_{X}}-\frac{\mu_{Y}}{\sigma_{Y}}\bigg{)}^{2j}\bigg{(}\frac{1-\rho}{1+\rho}\bigg{)}^{k-j}$
@@
The PDF $f_{S_{n}}(x)$ is bounded for all $x\in\mathbb{R}$ if and only if $n\geq 2$.
@@
It was shown in [14] that the PDF of $Z=_{d}S_{1}$ (here $=_{d}$ denotes equality in distribution) is unimodal with a logaritmic singularity at the mode $x=0$ (this can also be inferred from formula (2.6) by using the limiting forms (A.32) and (A.33)). For $n\geq 2$, the PDF (2.6) of $S_{n}$ is, however, bounded for all $x\in\mathbb{R}$, as can be seen from applying the limiting form to (A.33) to determine the behaviour as $x\rightarrow 0$, and due to the fact that the function $U(a,b,x)$ only has (possible) singularities in the limits $x\rightarrow 0$ and $x\rightarrow\pm\infty$, and the PDF is clearly bounded in the limits $x\rightarrow\pm\infty$ (and the asymptotic behaviour of the PDF in the limits $x\rightarrow\pm\infty$ is in fact given in Theorem 3.1 of [15], which confirms that $f_{S_{n}}(x)\rightarrow 0$ as $x\rightarrow\pm\infty$). ∎
@@
$f_{\overline{Z}_{n}}(x)$ $=\frac{n(1-\rho^{2})^{n/2-1}}{2^{n-1}\sigma_{X}\sigma_{Y}}\exp\bigg{(}-\frac{n}{2(1-\rho^{2})}\bigg{(}\frac{\mu_{X}^{2}}{\sigma_{X}^{2}}+\frac{\mu_{Y}^{2}}{\sigma_{Y}^{2}}-\frac{2\rho\mu_{X}\mu_{Y}}{\sigma_{X}\sigma_{Y}}\bigg{)}-\frac{n|x|}{\sigma_{X}\sigma_{Y}(1+\rho\,\mathrm{sgn}(x))}\bigg{)}$
×$\quad\times\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/8)^{k}}{k!\Gamma(n/2+a_{j,k}(x))}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\bigg{)}^{j}\bigg{(}\frac{\mu_{X}}{\sigma_{X}}-\frac{\mu_{Y}}{\sigma_{Y}}\bigg{)}^{2j}\bigg{(}\frac{1-\rho}{1+\rho}\bigg{)}^{k-j}$
@@
$f_{S_{n}}(x)$ $=\frac{(1-\rho^{2})^{n/2-1}}{2^{n-1}\sigma_{X}\sigma_{Y}}\exp\bigg{(}-\frac{n\mu_{X}^{2}}{\sigma_{X}^{2}(1+\rho)}-\frac{|x|}{\sigma_{X}\sigma_{Y}(1+\rho\,\mathrm{sgn}(x))}\bigg{)}\sum_{k=0}^{\infty}\frac{(n/2)^{k}}{k!\Gamma(n/2+a_{0,k}(x))}$
@@
$f_{S_{n}}(x)$ $=\frac{(1-\rho^{2})^{n/2-1}}{2^{n-1}\sigma_{X}\sigma_{Y}}\exp\bigg{(}-\frac{n\mu_{X}^{2}}{\sigma_{X}^{2}(1-\rho)}-\frac{|x|}{\sigma_{X}\sigma_{Y}(1+\rho\,\mathrm{sgn}(x))}\bigg{)}\sum_{k=0}^{\infty}\frac{(n/2)^{k}}{k!\Gamma(n/2+a_{k,k}(x))}$
@@
We therefore see that formula (2.6) reduces to a single infinite series in the cases $\mu_{X}/\sigma_{X}=\mu_{Y}/\sigma_{Y}$ and $\mu_{X}/\sigma_{X}=-\mu_{Y}/\sigma_{Y}$. This is in contrast to the formula (1) of [9] that reduces to a single infinite series if $\mu_{X}/\sigma_{X}=\rho\mu_{Y}/\sigma_{Y}$ or $\mu_{Y}/\sigma_{Y}=\rho\mu_{X}/\sigma_{X}$ (see (1.3)).
@@
Suppose $\rho=0$, $\mu_{Y}=0$, and $\mu_{X}\in\mathbb{R}$. Then, for $x\in\mathbb{R}$,
@@
Theorem 2.6 extends in a natural way the PDF $f_{Z}(x)$ described in (1.2) to the PDF $f_{S_{n}}(x)$. On calculating $(2j)!=\Gamma(2j+1)$ using the gamma duplication formula $\Gamma(2x)=\pi^{-1/2}2^{2x-1}\Gamma(x)\Gamma(x+1/2)$ see ([25, Section 5.5(iii)]) we see that formula (2.7) with $n=1$ indeed reduces to (1.2).
@@
$\frac{-(\mu_{X}^{2}+\mu_{Y}^{2}-2\rho\mu_{X}\mu_{Y})t^{2}+2\mu_{X}\mu_{Y}\mathrm{i}t}{2[1-(1+\rho)\mathrm{i}t][1+(1-\rho)\mathrm{i}t]}$
$=-\frac{(\mu_{X}^{2}+\mu_{Y}^{2}-2\rho\mu_{X}\mu_{Y})}{2(1-\rho^{2})}+\frac{1}{4(1-\rho^{2})}\bigg{(}\frac{(1+\rho)(\mu_{X}-\mu_{Y})^{2}}{(1+(1-\rho)\mathrm{i}t)}+\frac{(1-\rho)(\mu_{X}+\mu_{Y})^{2}}{(1-(1+\rho)\mathrm{i}t)}\bigg{)}.$
@@
×$\quad\times\exp\bigg{(}\frac{n}{4(1-\rho^{2})}\bigg{(}\frac{(1+\rho)(\mu_{X}-\mu_{Y})^{2}}{(1+(1-\rho)\mathrm{i}t)}+\frac{(1-\rho)(\mu_{X}+\mu_{Y})^{2}}{(1-(1+\rho)\mathrm{i}t)}\bigg{)}\bigg{)}$
@@
×$\quad\times\sum_{k=0}^{\infty}\frac{(n/4)^{k}}{k!(1-\rho^{2})^{k}}\bigg{(}\frac{(1+\rho)(\mu_{X}-\mu_{Y})^{2}}{(1+(1-\rho)\mathrm{i}t)}+\frac{(1-\rho)(\mu_{X}+\mu_{Y})^{2}}{(1-(1+\rho)\mathrm{i}t)}\bigg{)}^{k}$
$=\exp\bigg{(}-\frac{n(\mu_{X}^{2}+\mu_{Y}^{2}-2\rho\mu_{X}\mu_{Y})}{2(1-\rho^{2})}\bigg{)}\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/4)^{k}}{k!(1-\rho^{2})^{k}}\binom{k}{j}(1+\rho)^{j}(\mu_{X}-\mu_{Y})^{2j}$
@@
$f_{S_{n}}(x)$ $=\frac{1}{2\pi}\exp\bigg{(}-\frac{n(\mu_{X}^{2}+\mu_{Y}^{2}-2\rho\mu_{X}\mu_{Y})}{2(1-\rho^{2})}\bigg{)}\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/4)^{k}}{k!(1-\rho^{2})^{k}}\binom{k}{j}(1+\rho)^{j}(\mu_{X}-\mu_{Y})^{2j}$
@@
$=\frac{1}{2\pi}\exp\bigg{(}-\frac{n(\mu_{X}^{2}+\mu_{Y}^{2}-2\rho\mu_{X}\mu_{Y})}{2(1-\rho^{2})}\bigg{)}\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/4)^{k}}{k!(1-\rho^{2})^{n/2+k}}\binom{k}{j}$
@@
$f_{S_{n}}(x)$ $=\frac{1}{2\pi}\exp\bigg{(}-\frac{n(\mu_{X}^{2}+\mu_{Y}^{2}-2\rho\mu_{X}\mu_{Y})}{2(1-\rho^{2})}\bigg{)}\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/4)^{k}}{k!(1-\rho^{2})^{n/2+k}}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\bigg{)}^{j}$
@@
×$\quad\times\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/8)^{k}}{k!\Gamma(n/2+k-j)}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\bigg{)}^{j}(\mu_{X}-\mu_{Y})^{2j}\bigg{(}\frac{1-\rho}{1+\rho}\bigg{)}^{k-j}$
@@
$\quad\sum_{k=0}^{\infty}\sum_{j=0}^{k}\frac{(n/8)^{k}}{k!\Gamma(n/2+j)}\binom{k}{j}\bigg{(}\frac{1+\rho}{1-\rho}\bigg{)}^{j}(\mu_{X}-\mu_{Y})^{2j}\bigg{(}\frac{1-\rho}{1+\rho}\bigg{)}^{k-j}$
@@
$=\exp\bigg{(}-\frac{n\mu^{2}}{2}\bigg{)}\sum_{k=0}^{\infty}\frac{1}{k!}\bigg{(}\frac{n\mu^{2}}{2}\bigg{)}^{k}\frac{1}{(1+t^{2})^{n/2+k}}.$
@@
$f_{S_{n}}(x)$ $=\frac{1}{2\pi}\exp\bigg{(}-\frac{n\mu^{2}}{2}\bigg{)}\sum_{k=0}^{\infty}\frac{1}{k!}\bigg{(}\frac{n\mu^{2}}{2}\bigg{)}^{k}\int_{-\infty}^{\infty}\frac{\mathrm{e}^{-\mathrm{i}xt}}{(1+t^{2})^{n/2+k}}\,\mathrm{d}t$